OpenStudy (anonymous):

Medal and fan! Please help (:

OpenStudy (paki):

please post your question....

OpenStudy (anonymous):

It's the attachment (:

OpenStudy (paki):

no attachment's here... please re upload...

OpenStudy (anonymous):

what happened to the attachment?

OpenStudy (anonymous):

which Q are you after?

OpenStudy (anonymous):

Whichever one you can explain to me, besides the first one (:

OpenStudy (anonymous):

@paki do you want to help with 9 and 10 and I'll help with 11 and 12?

OpenStudy (paki):

just explain one... and clear his/her concept..... i think that will be good.... @kx2bay

OpenStudy (anonymous):

ok, in 11 you see that x ia double y, so y = x/2, right?

OpenStudy (anonymous):

Ohh, right (:

OpenStudy (paki):

nice :)

OpenStudy (anonymous):

with 12 I think they want you to write this f(x) = x + 1

OpenStudy (anonymous):

Alright, I'll see! Thank you so much guys (:

OpenStudy (anonymous):

I think 9 is y = x + 3

OpenStudy (anonymous):

Yeah, looks like (: thank you!

OpenStudy (anonymous):

with Q 10, best draw a graph if its not a straight line then we're stuck, then formula for straight line is y = mx + c, do you know what m and c are?

OpenStudy (anonymous):

m = y2-y1 ------ this is the slope, c is the y intercept x2-x1

OpenStudy (anonymous):

Uhmm isn't m for slope and c for y intercept? Sorry, bad at math :(

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Oh! Haha yay (:

OpenStudy (anonymous):

take any 2 points from the table for y2, y1, x2, x1, easy isnt it?

OpenStudy (anonymous):

Yees!

OpenStudy (anonymous):

example for m http://openstudy.com/study#/updates/53884986e4b0db48f0cec469

OpenStudy (paki):

nice explanation.... good... @kx2bay

OpenStudy (anonymous):

thanks guys

OpenStudy (paki):

pleasure :)

OpenStudy (anonymous):

Thanks guys!! Haha